Consider a firm that uses capital and labor as inputs and sells 5,000 units of output per year at the prevailing market price s10. also assumes that the total labor cost for the firm is s45,000 per year. further assume that the company's total capital stock is currently worth 100,000, that the return available to investors with comparable risk is 10 percent per year, and that there is no depreciation. is this a profitable company? explain your answer
1
Expert's answer
2021-10-26T17:23:19-0400

Step 1

Part (a) Explicit cost is the cost of factor of production in which the company actually pays. Here Explicit cost is just the cost of labour and cost of capital.

Explicit cost = 45000 + 100,000 = 145000

Step 2

Part b)

Implicit cost is the opportunity cost equal to what a firm must give up in order to use a factor of production. For eg , money given to run and setup a business.

Here implicit cost = return on capital by investing it some other place = 10, 000

Step 3

Part c)

Total revenue of the firm is

TR=$(5,000x10)=$50,000

Total cost of the production is

TC = variable cost + fixed cost

= 45000 + 1,00,000 = $145000

Accounting loss =$ (95000)

Step 4

Part d)

Economic profit/ loss

= TR - TC ( including opportunity cost )

= 50,000- 155,000 = $(105000)